OpenStudy (anonymous):

find the inverse of f informally. verify that f(f^-1(x))=x and f^-1(f(x))=x f(x)= (x-1)/5

OpenStudy (anonymous):

f(x)= (x-1)/5 ..............(1) let f(x)=y means x=f^-1(y) y= (x-1)/5 x=5y+1 hence f^-1(y) = 5y+1 replace y by x f^-1(x) = 5x+1...................(2) now from (1)and(2) f(f^-1(x))=f(5x+1)=(5x+1-1)/5 =x hence f(f^-1(x))=x f^-1(f(x)=f^-1((x-1)/5)=5{(x-1)/5}+1=x-1+1=x hence f^-1(f(x))=x

OpenStudy (anonymous):

Any form of function when mapping it's own inverse function maps the object back to itself.

OpenStudy (anonymous):

For that function: \[f^{-1}(x)=5x+1\]:) Do you need the steps?

OpenStudy (anonymous):

no friend ...because problem says to do that..thats why i have chosed long path.